A group of students decide to react magnesium oxide with dilute nitric acid in a laboratory. It is decided to start with a mass of 2.0g of magnesium oxide in a conical flask. What volume of nitric acid, of concentration 2.5 mol dm-3, will be required to react with all the magnesium oxide in the conical flask?
1
Expert's answer
2019-05-24T02:21:54-0400

Firstly we must wright the reaction:

2HNO3 (dilute) + MgO = Mg(NO3)2 + H2O

Now we count out the amount of MgO

n = m/Mr = 2/40 = 0.05 mol

The amount of HNO3 :

n = 0.05 × 2 = 0.1 mol;

The volume of nitric acid is:

V = n/c = 0.1/2.5 = 0.04 dm3 = 0.04 l.